Radio wave radiation falls in the wavelength region of 10.0 to 1000 meters.  

What is the wavelength of radio wave radiation that has an energy of 5.27×10-31 kJ/photon?
Answer #1

Energy of a photon,

E = hc/

where,

E = energy of 1 photon = 5.27 x 10-31 kJ = 5.27 x 10-28 J    ( 1 kJ = 1000 J)

h = Planck's constant = 6.626 x 10-34 J.s

c = speed of light in vacuum = 3 x 108 m/s

= wavelength of radiation

Now, substituting these values in the above equation to get,

5.27 x 10-28 J = (6.626 x 10-34 J.s)(3 x 108 m/s)/

or, = (6.626 x 10-34 J.s)(3 x 108 m/s)/(5.27 x 10-28 J)

          = 377 m

Hence, the wavelength of radiation = 377 m